xercise 10-22 (Algorithmic) (LO. 8) Pedro, who is a single taxpayer, had AGI of $350,000 for 2017. He incurred the following expenses during the year: Medical expenses before 10%-of-AGI limitation $21,500 State and local income taxes $10,500 Real estate taxes $2,100 Home mortgage interest $17,500 Charitable contributions $4,200 Deductible investment interest expense $1,575 Compute the amount of Pedro's itemized deductions after any applicable reductions and/or limitations. Round your computations to the nearest dollar and use rounded amounts in subsequent calculations.
Answer #1

Step 1: Calculate Value of Total Itemized Deductions before Phaseout

The value of total itemized deductions before phaseout is calculated as below:

Value of Itemized Deductions before Phaseout = (Medical Expenses Subject to a Maximum of 10% of AGI) + State and Local Income Taxes + Real Estate Taxes + Home Mortgage Interest + Charitable Contributions + Investment Interest Expense

Substituting values in the above formula, we get,

Value of Itemized Deductions before Phaseout = (0) + 10,500 + 2,100 + 17,500 + 4,200 + 1,575 = $35,875

_____

Step 2: Calculate Value of Total Itemized Deductions Subject to Overall Limitation

The value of total itemized deduction subject to overall limitation is determined as follows:

Value of Itemized Deductions Subject to Overall Limitation = State and Local Income Taxes + Real Estate Taxes + Home Mortgage Interest + Charitable Contributions

Substituting values in the above formula, we get,

Value of Itemized Deductions Subject to Overall Limitation = 10,500 + 2,100 + 17,500 + 4,200 = $34,300

_____

Step 3: Calculate Value of Reduction/Limitation

The value of applicable reduction/limitation is the lower of the following amounts:

3% of AGI above the applicable threshold = 3%*(350,000 - 261,500) = $2,655 (the threshold limit for 2017 for individual taxpayer is $261,500)

or

80% of the amount of itemized deductions otherwise allowable for the tax year. = 34,300*80% = $27,440

Therefore, the amount of applicable reduction/limitation would be $2,655.

_____

Step 4: Calculate Pedro's Itemized Deductions after any Applicable Reductions and/or Limitations

The amount of Pedro's itemized deductions after any applicable reductions and/or limitations is calculated as below:

Pedro's Itemized Deductions after any Applicable Reductions and/or Limitations = Value of Itemized Deductions before Phaseout - Amount of Applicable Reduction/Limitation = 35,875 - 2,655 = $33,220 (answer)